# MetaRAG: Metamorphic Testing for Hallucination Detection in RAG Systems

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in enterprise applications, yet their reliability remains limited by hallucinations, i.e., confident but factually incorrect information. Existing detection approaches, such as SelfCheckGPT and MetaQA, primarily target standalone LLMs and do not address the unique challenges of Retrieval-Augmented Generation (RAG) systems, where responses must be consistent with retrieved evidence. We therefore present MetaRAG, a metamorphic testing framework for hallucination detection in Retrieval-Augmented Generation (RAG) systems. MetaRAG operates in a real-time, unsupervised, black-box setting, requiring neither ground-truth references nor access to model internals, making it suitable for proprietary and high-stakes domains. The framework proceeds in four stages: (1) decompose answers into atomic factoids, (2) generate controlled mutations of each factoid using synonym and antonym substitutions, (3) verify each variant against the retrieved context (synonyms are expected to be entailed and antonyms contradicted), and (4) aggregate penalties for inconsistencies into a response-level hallucination score. Crucially for identity-aware AI, MetaRAG localizes unsupported claims at the factoid span where they occur (e.g., pregnancy-specific precautions, LGBTQ+ refugee rights, or labor eligibility), allowing users to see flagged spans and enabling system designers to configure thresholds and guardrails for identity-sensitive queries. Experiments on a proprietary enterprise dataset illustrate the effectiveness of MetaRAG for detecting hallucinations and enabling trustworthy deployment of RAG-based conversational agents. We also outline a topic-based deployment design that translates MetaRAG’s span-level scores into identity-aware safeguards; this design is discussed but not evaluated in our experiments.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 and Llama-3 are transforming enterprise applications in healthcare, law, and customer service [8, 7, 2]. They power chatbots and virtual assistants that interact in natural language, offering unprecedented convenience and efficiency [30]. However, as these systems move into production, a persistent challenge emerges: hallucinations, i.e., responses that are fluent and convincing but factually incorrect or unsupported by evidence [23, 14].
In high-stakes domains such as healthcare, law, and finance, hallucinations are not merely a nuisance but a critical barrier to reliable adoption, raising concerns about user trust, regulatory compliance, and business risk [20]. Moreover, hallucinations are not uniformly risky: the same unsupported claim can differentially affect specific populations. In healthcare (e.g., pregnancy/trimester-specific contraindications), migration and asylum (e.g., protections for LGBTQ+ refugees), or labor rights (e.g., eligibility by status), ungrounded spans can cause disproportionate harm. Rather than treating users as homogeneous, hallucination detection methods should make such spans reviewable at the factoid level so downstream systems can apply identity-aware policies (e.g., stricter thresholds, forced citations, or escalation to a human) when the topic indicates elevated risk. This perspective connects hallucination detection to identity-aware deployment, where span-level evidence enables topic-conditioned safeguards that reduce disproportionate risk.
Ji et al. [14] categorize hallucinations into two types:
@@
Retrieval-Augmented Generation (RAG) [17] aims to mitigate hallucinations by grounding model outputs in retrieved, up-to-date documents, as illustrated in Figure 1. By injecting retrieved text from reliable external sources and proprietary documents, into the prompt, RAG improves factuality and domain relevance. While effective against intrinsic hallucinations, RAG remains susceptible to extrinsic hallucinations, especially when retrieved evidence is ignored, misinterpreted, or insufficient [10].
Detecting hallucinations is particularly challenging in real-world settings, where RAG-based chatbots must respond to queries about unseen, proprietary, or confidential content where gold-standard references are typically unavailable [19]. Many existing hallucination detection methods rely on gold-standard reference answers [23, 16], annotated datasets [33], or access to model internals such as hidden states or token log-probabilities [25, 6]. However, in enterprise settings, such internals are often inaccessible: many state-of-the-art LLMs (e.g., GPT-4, Claude) are proprietary and only accessible via APIs that expose the final output text but not intermediate computations, limiting the feasibility of these methods in practice [19].
@@
  • We analyze the performance–latency/cost trade-offs of MetaRAG and provide a consistency analysis to guide future research and practical deployment.
  • We outline identity-aware safeguards (topic-aware thresholds, forced citation, escalation) that consume MetaRAG’s scores; these safeguards are a deployment design and are not part of our empirical evaluation.
@@ 2.1 Definitions of Hallucination
The term hallucination has been used with varying scope across natural language generation tasks. Some studies emphasize factuality, describing hallucinations as outputs that contradict established facts, i.e., inconsistencies with world knowledge or external ground truth [31, 13]. Others highlight faithfulness, where hallucinations occur when generated responses deviate from the user instruction or a reference text, often producing plausible but ungrounded statements particularly in source-conditioned tasks such as summarization or question answering [26]. Beyond these two dimensions, researchers also note cases of incoherent or nonsensical text that cannot be clearly attributed to factuality or faithfulness criteria  [14, 23].
Alternative terms have also been introduced. Confabulation draws on psychology to describe fluent but fabricated content arising from model priors [9], while fabrication is preferred by some to avoid anthropomorphic connotations [1, 21]. More recently, Chakraborty et al. [3] propose a flexible definition tailored to deployment settings, defining a hallucination as a generated output that conflicts with constraints or deviates from desired behavior in actual deployment, while remaining syntactically plausible under the circumstance.
@@ 2.2 Hallucination Detection in LLMs
Building on these definitions, hallucinations have been recognized as a major challenge in text generation. Early work in machine translation and abstractive summarization described them as outputs that are not grounded in the input source [23, 16, 15], motivating the development of evaluation metrics and detection methods for faithfulness and factual consistency across natural language generation tasks.
More recent reference-free (unsupervised or zero-reference) methods aim to detect hallucinations without gold-standard labels by analyzing the model’s own outputs. A prominent method is SelfCheckGPT [22], a zero-resource, black-box approach that queries the LLM multiple times with the same prompt and measures semantic consistency across responses. The intuition is that hallucinated content often leads to instability under stochastic re-generation; true facts remain stable, while fabricated ones diverge. Manakul et al. show that SelfCheckGPT achieves strong performance in sentence-level hallucination detection compared to gray-box methods, and emphasize that it requires no external database or access to model internals [22]. However, SelfCheckGPT may struggle when deterministic decoding or high model confidence leads to repeating the same incorrect output.
@@ 2.3 Metamorphic Testing
Metamorphic Testing (MT) [4] was originally proposed in software engineering to address the oracle problem in which the correct output is unknown. MT relies on metamorphic relations (MRs): transformations of the input with predictable effects on outputs, enabling error detection without access to ground truth [28]. In machine learning, MT has been applied to validate models in computer vision  [5] (e.g., rotating an image should not change its predicted class) and NLP [27]
In hallucination detection for LLMs, MetaQA [32] leverages MRs by generating paraphrased or antonym-based question variants and verifying whether answers satisfy expected semantic or logical constraints. Relying purely on prompt mutations and consistency checks, MetaQA achieves higher precision and recall than SelfCheckGPT on open-domain QA.
Researchers have also adapted MT for more complex conversational and reasoning settings. MORTAR [12] applies dialogue-level perturbations and knowledge-graph-based inference to multi-turn systems, detecting up to four times more unique bugs than single-turn MT. Drowzee [18] uses logic programming to construct temporal and logical rules from Wikipedia, generating fact-conflicting test cases and revealing rates of 24.7% to 59.8% across six LLMs in nine domains [32].
@@ 3.1 Overview
Building on the metamorphic testing (MT) methodology to detect hallucinations in LLMs introduced by MetaQA [32], MetaRAG advances this approach to detect hallucinations in retrieval-augmented generation (RAG) settings by introducing a context-based verification stage. A metamorphic testing layer operates on top of the standard RAG pipeline to automatically detect hallucinated responses. Figure 2 outlines the workflow.
@@
Figure 2: Overview of the MetaRAG workflow. (A) Integration of MetaRAG with a standard RAG pipeline: given a user question, the RAG retrieves context and generates an answer, which is then passed to MetaRAG for hallucination detection. (B) Internal MetaRAG pipeline: the answer is decomposed into atomic factoids, each factoid is mutated through synonym and antonym substitutions, and verified against the retrieved context using entailment/contradiction checks. Penalties are assigned to inconsistencies, and scores are aggregated into a response-level hallucination score.
@@ 3.3 Step 2: Mutation Generation
Each factoid (hereafter, fact) from Step 1, MetaRAG applies metamorphic mutations to generate perturbed variants of the original claim. This step is grounded in the principle of metamorphic testing, where controlled semantic transformations are used to probe model consistency and expose hallucinations [32].
@@ 5.2 Results
Table 2: Validation leaderboards for 26 MetaRAG configurations, showing the top four for each of F1 score, Precision, and Recall.
Top–4 by ID Config. F1 Prec. Rec. Acc.
F1 5 mini/41/multi/2/0 0.9391 1.0000 0.8853 0.9401
18 mini/mini/41/5/0.7 0.9372 0.9087 0.9676 0.9401
19 mini/mini/41/5/0 0.9352 0.9352 0.9352 0.9400
16 mini/41/multi/5/0 0.9228 0.8819 0.9676 0.9250
Precision 22 mini/mini/multi/5/0 0.9149 0.9641 0.8704 0.9250
19 mini/mini/41/5/0 0.9352 0.9352 0.9352 0.9400
18 mini/mini/41/5/0.7 0.9372 0.9087 0.9676 0.9401
24 41/mini/multi/5/0 0.8847 0.8996 0.8704 0.8951
Recall 1 mini/41/41/2/0.7 0.8736 0.7755 1.0000 0.8660
5 mini/41/multi/2/0 0.9391 0.8853 1.0000 0.9401
9 mini/mini/mini/2/0.7 0.9114 0.8372 1.0000 0.9101
4 mini/41/41/2/0 0.8614 0.7565 1.0000 0.8510
@@ 6.2 Ethical Considerations
  • Reinforced stereotypes: Language models are known to reproduce and amplify societal biases, as demonstrated by benchmarks such as StereoSet [24] and WinoBias [34]. In identity-sensitive deployments, hallucinated outputs risk reinforcing these biases in subtle but harmful ways.
  • Trust erosion: Chatbots are only adopted at scale in high-stakes domains if users trust their outputs. Surveys on hallucination consistently highlight that exposure to unsupported or fabricated content undermines user trust in LLM systems [14, 20].
  • Identity harms: Misrepresentations in generated responses may distort personal narratives or marginalize underrepresented groups, aligning with broader critiques that technical systems can reproduce social inequities if identity considerations are overlooked [11, 29].
By detecting hallucinations in a black-box, reference-free manner, MetaRAG supports safer deployment of RAG-based systems, particularly in settings where fairness, identity, and user well-being are at stake.
@@ 6.3 Limitations and Future Work
  • Dataset Scope: The study relies on a private, domain-specific dataset. This may limit external validity. Future work should focus on curating or constructing public benchmarks designed to avoid overlap with LLM pretraining corpora, enabling more robust generalizability.

### v2

# Identity-Aware AI workshop at 28th European Conference on Artificial Intelligence, October 25, 2025, Bologna, Italy

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in enterprise applications, yet their reliability remains limited by hallucinations, i.e., confident but factually incorrect information. Existing detection approaches, such as SelfCheckGPT and MetaQA, primarily target standalone LLMs and do not address the unique challenges of Retrieval-Augmented Generation (RAG) systems, where responses must be consistent with retrieved evidence. We therefore present MetaRAG, a metamorphic testing framework for hallucination detection in Retrieval-Augmented Generation (RAG) systems. MetaRAG operates in a real-time, unsupervised, black-box setting, requiring neither ground-truth references nor access to model internals, making it suitable for proprietary and enterprise deployments. The framework proceeds in four stages: (1) decompose answers into atomic factoids, (2) generate controlled mutations of each factoid using synonym and antonym substitutions, (3) verify each variant against the retrieved context (synonyms are expected to be entailed and antonyms contradicted), and (4) aggregate penalties for inconsistencies into a response-level hallucination score. MetaRAG further localizes unsupported claims at the span level, enabling transparent visualization of potentially hallucinated segments and supporting configurable safeguards in sensitive use cases. Experiments on a proprietary enterprise dataset demonstrate the effectiveness of MetaRAG for detecting hallucinations and enabling trustworthy deployment of RAG-based conversational agents. We also outline a topic-based deployment design that translates MetaRAG’s span-level scores into identity-aware safeguards; this design is discussed but not evaluated in our experiments.
keywords:
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 and Llama-3 are transforming enterprise applications in healthcare, law, and customer service [openai2024gpt4technicalreport, meta2024llama3, bommasani2021opportunities]. They power chatbots and virtual assistants that interact in natural language, offering unprecedented convenience and efficiency [singhal2023medpalm]. However, as these systems move into production, a persistent challenge emerges: hallucinations, i.e., responses that are fluent and convincing but factually incorrect or unsupported by evidence [maynez-etal-2020-faithfulness, 10.1145/3571730survey].
In domains such as healthcare, law, and finance, hallucinations are not merely a nuisance but a critical barrier to reliable adoption, raising concerns about user trust, regulatory compliance, and business risk [lyu2024trustworthy]. Moreover, hallucinations are not uniformly risky: the same unsupported claim can differentially affect specific populations. In healthcare (e.g., pregnancy/trimester-specific contraindications), migration and asylum (e.g., protections for LGBTQ+ refugees), or labor rights (e.g., eligibility by status), ungrounded spans can cause disproportionate harm. Rather than treating users as homogeneous, hallucination detection methods should make such spans reviewable at the factoid level so downstream systems can apply identity-aware policies (e.g., stricter thresholds, forced citations, or escalation to a human) when the topic indicates elevated risk. This perspective connects hallucination detection to identity-aware deployment, where span-level evidence enables topic-conditioned safeguards that reduce disproportionate risk.
Ji et al. [10.1145/3571730survey] categorize hallucinations into two types:
@@
Retrieval-Augmented Generation (RAG) [lewis2020rag] aims to mitigate hallucinations by grounding model outputs in retrieved, up-to-date documents, as illustrated in Figure 1. By injecting retrieved text from reliable external sources and proprietary documents, into the prompt, RAG improves factuality and domain relevance. While effective against intrinsic hallucinations, RAG remains susceptible to extrinsic hallucinations, especially when retrieved evidence is ignored, misinterpreted, or insufficient [gao2023faithfulrag].
Detecting hallucinations is particularly challenging in real-world settings, where RAG-based chatbots must respond to queries about unseen, proprietary, or confidential content where gold-standard references are typically unavailable [liang2022holistic]. Many existing hallucination detection methods rely on gold-standard reference answers [maynez-etal-2020-faithfulness, kryscinski-etal-2020-evaluating], annotated datasets [zhang-etal-2024-benchmarking], or access to model internals such as hidden states or token log-probabilities [rashkin-etal-2023-measuring, dziri-etal-2022-faithdial]. However, in enterprise settings, such internals are often inaccessible: many state-of-the-art LLMs (e.g., GPT-4, Claude) are proprietary and only accessible via APIs that expose the final output text but not intermediate computations, limiting the feasibility of these methods in practice [liang2022holistic].
@@
  • We analyze the performance–latency/cost trade‑offs of MetaRAG and provide a consistency analysis to guide future research and practical deployment.
  • We outline identity-aware safeguards (topic-aware thresholds, forced citation, escalation) that can consume MetaRAG’s scores; these safeguards are a deployment design and are not part of our empirical evaluation.
@@ 2.1 Definitions of Hallucination
The term hallucination has been used with varying scope across natural language generation tasks. Some studies emphasize factuality, describing hallucinations as outputs that contradict established facts, i.e., inconsistencies with world knowledge or external ground truth [wang-etal-2024-factuality, Huang_2025]. Others highlight faithfulness, where hallucinations occur when generated responses deviate from the user instruction or a reference text, often producing plausible but ungrounded statements particularly in source-conditioned tasks such as summarization or question answering [rawte2023surveyhallucinationlargefoundation]. Beyond these two dimensions, researchers also note cases of incoherent or nonsensical text that cannot be clearly attributed to factuality or faithfulness criteria  [10.1145/3571730survey, maynez-etal-2020-faithfulness].
Alternative terms have also been introduced. Confabulation draws on psychology to describe fluent but fabricated content arising from model priors [sui-etal-2024-confabulation], while fabrication is preferred by some to avoid anthropomorphic connotations [azamfirei2023hallucinations, maleki2024aihallucinationsmisnomerworth]. More recently, Chakraborty et al. [chakraborty2025] propose a flexible definition tailored to deployment settings, defining a hallucination as a generated output that conflicts with constraints or deviates from desired behavior in actual deployment, while remaining syntactically plausible under the circumstance.
@@ 2.2 Hallucination Detection in LLMs
Building on these definitions, hallucinations have been recognized as a major challenge in text generation. Early work in machine translation and abstractive summarization described them as outputs that are not grounded in the input source [maynez-etal-2020-faithfulness, kryscinski-etal-2020-evaluating, koehn-knowles-2017-six], motivating the development of evaluation metrics and detection methods for faithfulness and factual consistency across natural language generation tasks.
More recent reference-free (unsupervised or zero-reference) methods aim to detect hallucinations without gold-standard labels by analyzing the model’s own outputs. A prominent method is SelfCheckGPT [manakul2023selfcheckgptzeroresourceblackboxhallucination], a zero-resource, black-box approach that queries the LLM multiple times with the same prompt and measures semantic consistency across responses. The intuition is that hallucinated content often leads to instability under stochastic re-generation; true facts remain stable, while fabricated ones diverge. Manakul et al. show that SelfCheckGPT achieves strong performance in sentence-level hallucination detection compared to gray-box methods, and emphasize that it requires no external database or access to model internals [manakul2023selfcheckgptzeroresourceblackboxhallucination]. However, SelfCheckGPT may struggle when deterministic decoding or high model confidence leads to repeating the same incorrect output.
@@ 2.3 Metamorphic Testing
Metamorphic Testing (MT) [chen2020metamorphictestingnewapproach] was originally proposed in software engineering to address the oracle problem in which the correct output is unknown. MT relies on metamorphic relations (MRs): transformations of the input with predictable effects on outputs, enabling error detection without access to ground truth [7422146]. In machine learning, MT has been applied to validate models in computer vision  [Dwarakanath_2018] (e.g., rotating an image should not change its predicted class) and NLP [ribeiro-etal-2020-beyond]
In hallucination detection for LLMs, MetaQA [yang2025hallucinationdetectionlargelanguage] leverages MRs by generating paraphrased or antonym-based question variants and verifying whether answers satisfy expected semantic or logical constraints. Relying purely on prompt mutations and consistency checks, MetaQA achieves higher precision and recall than SelfCheckGPT on open-domain QA.
Researchers have also adapted MT for more complex conversational and reasoning settings. MORTAR [guo2025mortarmultiturnmetamorphictesting] applies dialogue-level perturbations and knowledge-graph-based inference to multi-turn systems, detecting up to four times more unique bugs than single-turn MT. Drowzee [li2024drowzeemetamorphictestingfactconflicting] uses logic programming to construct temporal and logical rules from Wikipedia, generating fact-conflicting test cases and revealing rates of 24.7% to 59.8% across six LLMs in nine domains [yang2025hallucinationdetectionlargelanguage].
@@ 3.1 Overview
Building on the metamorphic testing (MT) methodology to detect hallucinations in LLMs introduced by MetaQA [yang2025hallucinationdetectionlargelanguage], MetaRAG advances this approach to detect hallucinations in retrieval-augmented generation (RAG) settings by introducing a context-based verification stage. A metamorphic testing layer operates on top of the standard RAG pipeline to automatically detect hallucinated responses. Figure 2 outlines the workflow.
@@
Figure 2: Overview of the MetaRAG workflow. (A) Integration of MetaRAG into a standard RAG pipeline: given a user question, the RAG retrieves context and generates an answer, which is then passed to MetaRAG for hallucination detection. (B) Internal MetaRAG pipeline: the answer is decomposed into atomic factoids, each factoid is mutated through synonym and antonym substitutions, and verified against the retrieved context using entailment/contradiction checks. Penalties are assigned to inconsistencies, and scores are aggregated into a response-level hallucination score.
@@ 3.3 Step 2: Mutation Generation
Each factoid (hereafter, fact) from Step 1, MetaRAG applies metamorphic mutations to generate perturbed variants of the original claim. This step is grounded in the principle of metamorphic testing, where controlled semantic transformations are used to probe model consistency and expose hallucinations [yang2025hallucinationdetectionlargelanguage].
@@ 5.2 Results
Table 2: Validation leaderboards for 26 MetaRAG configurations, showing the top–4 for each metric.
ID Config. F1 Prec. Rec. Acc.
Top-4 by F1
5 mini/41/multi/2/0 0.939 1.000 0.885 0.940
18 mini/mini/41/5/0.7 0.937 0.909 0.968 0.940
19 mini/mini/41/5/0 0.935 0.935 0.935 0.940
16 mini/41/multi/5/0 0.923 0.882 0.968 0.925
Top-4 by Precision
22 mini/mini/multi/5/0 0.915 0.964 0.870 0.925
19 mini/mini/41/5/0 0.935 0.935 0.935 0.940
18 mini/mini/41/5/0.7 0.937 0.909 0.968 0.940
24 41/mini/multi/5/0 0.885 0.900 0.870 0.895
Top-4 by Recall
1 mini/41/41/2/0.7 0.874 0.776 1.000 0.866
5 mini/41/multi/2/0 0.939 0.885 1.000 0.940
9 mini/mini/mini/2/0.7 0.911 0.837 1.000 0.910
4 mini/41/41/2/0 0.861 0.757 1.000 0.851
@@ 6.2 Ethical Considerations
  • Reinforced stereotypes: Language models are known to reproduce and amplify societal biases, as demonstrated by benchmarks such as StereoSet [nadeem-etal-2021-stereoset] and WinoBias [zhao-etal-2018-gender]. In identity-sensitive deployments, hallucinated outputs risk reinforcing these biases in subtle but harmful ways.
  • Trust erosion: Chatbots are only adopted at scale in high-stakes domains if users trust their outputs. Surveys on hallucination consistently highlight that exposure to unsupported or fabricated content undermines user trust in LLM systems [10.1145/3571730survey, lyu2024trustworthy].
  • Identity harms: Misrepresentations in generated responses may distort personal narratives or marginalize underrepresented groups, aligning with broader critiques that technical systems can reproduce social inequities if identity considerations are overlooked [gebru2021datasheets, selbst2019fairness].
By detecting hallucinations in a black-box, reference-free manner, MetaRAG can support safer deployment of RAG-based systems, particularly in settings where fairness, identity, and user well-being are at stake.
@@ 6.3 Limitations and Future Work
  • Dataset Scope: The study relies on a private, domain-specific dataset. This may limit external validity. Future work should focus on curating or constructing public benchmarks designed to avoid overlap with LLM pretraining corpora, enabling more robust generalization.